# The theory of implicit operations

### v1

@@ 4. Existential elimination
As $\bm{A}\leqslant\bm{B}\leqslant\bm{T}_{\mathsf{K}}(\varphi)\in\mathsf{K}$ and $\mathsf{K}$ is a quasivariety with the amalgamation property, we can apply Proposition 4.11 to $y/\theta(\varphi)\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{T}_{\mathsf{K}}(\varphi))\cap B$, obtaining $y/\theta(\varphi)\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. By Theorem 4.10 there exist an $m$-ary $g\in\mathsf{imp}_{pp}(\mathsf{K})$ and $\langle a_{1},\dots,a_{m}\rangle\in\mathsf{dom}(g^{\bm{B}})\cap A^{m}$ such that $g^{\bm{B}}(a_{1},\dots,a_{m})=y/\theta(\varphi)$.
Since $g\in\mathsf{imp}_{pp}(\mathsf{K})$, there exists a formula $\exists z_{1},\dots,z_{k}\psi(x_{1},\dots,x_{m},y,z_{1},\dots,z_{k})$, where $\psi$ is a conjunction of equations, defining $g$. Together with $\langle a_{1},\dots,a_{m}\rangle\in\mathsf{dom}(g^{\bm{B}})$ and $g^{\bm{B}}(a_{1},\dots,a_{n})=y/\theta(\varphi)$, this guarantees the existence of $b_{1},\dots,b_{k}\in B$ such that
@@
Notice that $\gamma$ is a conjunction of equations because so is $\psi$. Then observe that the formula $\exists z_{1},\dots,z_{k}\psi(x_{1},\dots,x_{m},y,z_{1},\dots,z_{k})$ is functional in $\mathsf{K}$ because it defines $g$. Together with the definition of $\gamma$, this guarantees that $\gamma$ is also functional in $\mathsf{K}$. Hence, $\gamma$ defines some $h\in\mathsf{imp}_{eq}(\mathsf{K})$ by Corollary 3.11. Therefore, to conclude the proof, it suffices to show that $h$ interpolates $f$.
@@
$\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=\{b\in B:$ $\text{ there exist }f\in\mathsf{imp}_{eq}(\mathsf{K})\text{ and }\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$
@@
We close this section with two examples. The first shows that Isbell’s operations cannot be interpolated by any implicit operation defined by a conjunction of equations in the variety of commutative monoids, and the second shows that condition (i) of Theorem 4.3 cannot be improved by requiring the equality $\mathsf{imp}_{pp}(\mathsf{K})=\mathsf{imp}_{eq}(\mathsf{K})$.
@@
For every positive $n$ let $f_{n}$ be the $n$-th Isbell’s operation, viewed as an implicit operation of the variety $\mathsf{Mon}$ of monoids (see 3.14 and 3.15). Then $f_{n}\in\mathsf{imp}_{pp}(\mathsf{Mon})$. We will prove that $f_{n}$ cannot be interpolated by any member of $\mathsf{imp}_{eq}(\mathsf{\mathsf{Mon}})$. Suppose the contrary, with a view to contradiction. Then $f_{n}$ is interpolated by some $g\in\mathsf{imp}_{eq}(\mathsf{Mon})$. Consider the commutative monoids $\mathbb{N}=\langle\mathbb{N};\cdot,1\rangle$ and $\mathbb{Q}=\langle\mathbb{Q};\cdot,1\rangle$. Moreover, let $a_{1},\dots,a_{n},b_{1},\dots,b_{n},c_{1},\dots,c_{2n+1},d$ be the sequence of rationals defined as follows: for every $1\leqslant i\leqslant n$ and $1<j<2n+1$,
@@
Using the formula $\psi_{n}$ in 3.14, we have
@@
By the definition of $f_{n}$ this yields $\langle c_{1},\dots,c_{2n+1}\rangle\in\mathsf{dom}(f_{n}^{\mathbb{Q}})$ and $f_{n}^{\mathbb{Q}}(c_{1},\dots,c_{2n+1})=d$. As $g$ interpolates $f_{n}$, we obtain $\langle c_{1},\dots,c_{2n+1}\rangle\in\mathsf{dom}(g^{\mathbb{Q}})$ and $g^{\mathbb{Q}}(c_{1},\dots,c_{2n+1})=d$. Let $\varphi(x_{1},\dots,x_{2n+1},y)$ be the conjunction of equations defining $g$. Then $\mathbb{Q}\vDash\varphi(c_{1},\dots,c_{2n+1},d)$. As $\varphi$ is a universal formula and $c_{1},\dots,c_{2n+1},d\in\mathbb{N}$, 1.1(iii) implies that $\mathbb{N}\vDash\varphi(c_{1},\dots,c_{2n+1},d)$, and hence $\langle c_{1},\dots,c_{2n+1}\rangle\in\mathsf{dom}(g^{\mathbb{N}})$ and $g^{\mathbb{N}}(c_{1},\dots,c_{2n+1})=d$. Let $\bm{A}$ and $\bm{B}$ be the submonoids of $\mathbb{N}$ with universes $A=\{1\}\cup\{2m:m\in\mathbb{N}\}$ and $B=\{0,1\}$, respectively. Since $c_{1},\dots,c_{2n+1}\in A$ and $g\in\mathsf{imp}_{eq}(\mathsf{Mon})$, 4.10 yields $3=d\in\mathsf{d}_{\mathsf{Mon}}(\bm{A},\mathbb{N})$. Let $h,k\colon\mathbb{N}\to\bm{B}$ be given by
@@
It is immediate to verify that $h$ and $k$ are homomorphisms such that $h{\upharpoonright}_{A}=k{\upharpoonright}_{A}$ and $h(3)\neq k(3)$, whence $3\notin\mathsf{d}_{\mathsf{Mon}}(\bm{A},\mathbb{N})$. As this is false, we conclude that $f_{n}$ cannot be interpolated by any member of $\mathsf{imp}_{eq}(\mathsf{Mon})$. ∎
@@
We recall that $\mathsf{CCMon}$ has the amalgamation property (see, e.g., [76, pp. 100, 108]) and, therefore, falls within the scope of 4.3. On the other hand, we will show that $\mathsf{imp}_{pp}(\mathsf{CCMon})\neq\mathsf{imp}_{eq}(\mathsf{CCMon})$.
@@
Notice that $\varphi$ is functional in $\mathsf{CCMon}$. Indeed, $\bm{A}\vDash\varphi(a,b)$ implies $b=1$ for all $\bm{A}\in\mathsf{CCMon}$ and $a,b\in A$. Since $\varphi$ is a pp formula, we can apply 3.11, obtaining that it defines a unary $f\in\mathsf{imp}_{pp}(\mathsf{CCMon})$. Moreover, if $\bm{A}\in\mathsf{CCMon}$, then $\mathsf{dom}(f^{\bm{A}})$ consists of the invertible elements of $\bm{A}$. We show that $f\notin\mathsf{imp}_{eq}(\mathsf{CCMon})$. Suppose the contrary, with a view to contradiction. Then consider the cancellative commutative monoids $\mathbb{N}=\langle\mathbb{N};\cdot,1\rangle$ and $\mathbb{Q}=\langle\mathbb{Q};\cdot,1\rangle$. We have
@@
Since $f$ is defined by a conjunction of equations, from the above display and $1,2\in\mathbb{N}\leqslant\mathbb{Q}$ it follows that $2\in\mathsf{dom}(f^{\mathbb{N}})$, a contradiction with the fact that $2$ is not invertible in $\mathbb{N}$. Hence, we conclude that $\mathsf{imp}_{pp}(\mathsf{CCMon})\neq\mathsf{imp}_{eq}(\mathsf{CCMon})$.
We conclude this example by showing that $\mathsf{imp}_{eq}(\mathsf{CCMon})$ is not closed under composition. Observe that $f$ coincides with the composition $h\circ g$, where $h\in\mathsf{imp}_{eq}(\mathsf{CCMon})$ is the unary implicit operation defined by the equation $y\thickapprox 1$ and $g\in\mathsf{imp}_{eq}(\mathsf{CCMon})$ is the implicit operation of “taking inverses” in monoids (see 3.4) restricted to $\mathsf{CCMon}$. Since $h\circ g=f\notin\mathsf{imp}_{eq}(\mathsf{CCMon})$, this shows that $\mathsf{imp}_{eq}(\mathsf{CCMon})$ is not closed under composition. ∎
@@ 5. The strong Beth definability property
The next definition formalizes the idea of interpolating implicit operations by sets of terms and is a particular instance of the notion of interpolation introduced in 4.2.
@@
As a consequence of the 1.3, we obtain the following.
@@
Let $f$ be a partial function on an elementary class $\mathsf{K}$. Assume that $f$ is defined by a formula $\varphi(x_{1},\dots,x_{n},y)$ and that it can be interpolated by a set of terms $\{t_{i}:i\in I\}$. Then condition (21) holds. From the 1.3 it follows that there exists a finite $T\subseteq\{t_{i}:i\in I\}$ such that
@@
When viewed as an implicit operation on a class of algebras $\mathsf{K}$, every term function $\langle t^{\bm{A}}:\bm{A}\in\mathsf{K}\rangle$ of $\mathsf{K}$ is interpolated by a single term, namely, $t$ (see Example 3.8). However, not all implicit operations can be interpolated by terms. For instance, there is no set of terms interpolating the implicit operation of “taking inverses” in the variety of monoids as we will see in 6.8. It is therefore sensible to isolate the cases in which interpolation is always possible, something that indicates a good balance between the expressivity of the language (measured by what can be said in terms of implicit operations) and its actual richness (measured by what can be interpolated, or made explicit, by terms).
@@ 6. The strong epimorphism surjectivity property
We will show that, in the setting of universal classes, the strong Beth definability property is equivalent to the following strengthening of the epimorphism surjectivity property (see 6.5).
@@
While every class with the strong epimorphism surjectivity property has the epimorphism surjectivity property, the converse need not hold in general. For instance, it is known that only $16$ varieties of Heyting algebras have the strong epimorphism surjectivity property [91, Thm. 8.1]. On the other hand, there exists a continuum of varieties of Heyting algebras with the nonstrong version of this property [14, p. 199]. However, the two properties coincide in quasivarieties with the amalgamation property (see 7.14).
@@
In the context of quasivarieties, the strong epimorphism surjectivity property admits a purely categorical formulation. For observe that each quasivariety $\mathsf{K}$ can be viewed as a category whose objects are the members of $\mathsf{K}$ and whose arrows are the homomorphisms between them. As quasivarieties contain free algebras (see 2.19), monomorphisms coincide with embeddings in quasivarieties (see, e.g., [1, Prop. 8.29]). A monomorphism is said to be regular when it is an equalizer. It turns out that a quasivariety $\mathsf{K}$ has the strong epimorphism surjectivity property if and only if all monomorphisms are regular in $\mathsf{K}$ (see, e.g., [76, Prop. 6.1]). ∎
@@
As an immediate consequence of 6.2, we obtain an alternative formulation of the strong epimorphism surjectivity property in terms of dominions.
@@
From 4.10 and 6.6 we deduce the following.
@@
A universal class $\mathsf{K}$ has the strong epimorphism surjectivity property if and only if for all $\bm{A}\leqslant\bm{B}\in\mathsf{K}$, $f\in\mathsf{imp}_{pp}(\mathsf{K})$, and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$ we have $f^{\bm{B}}(a_{1},\dots,a_{n})\in A$.
We are now ready to prove 6.5.
@@
Let $\mathsf{K}$ be a universal class. To prove the implication from left to right, we reason by contraposition. Suppose that $\mathsf{K}$ lacks the strong Beth definability property. By Propositions 5.2 and 5.4 there exists $f\in\mathsf{imp}_{pp}(\mathsf{K})$ that cannot be interpolated by any set of terms. Therefore, there exists $\bm{B}\in\mathsf{K}$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$ for which there exists no term $t$ such that $f^{\bm{B}}(a_{1},\dots,a_{n})=t^{\bm{B}}(a_{1},\dots,a_{n})$. Then $f^{\bm{B}}(a_{1},\dots,a_{n})\notin\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n})$. Let $\bm{A}=\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n})$. Since $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$ and $f^{\bm{B}}(a_{1},\dots a_{n})\notin A$, 6.7 implies that $\mathsf{K}$ lacks the strong epimorphism surjectivity property.
Next we prove the implication from right to left. Suppose that $\mathsf{K}$ has the strong Beth definability property. In order to prove that $\mathsf{K}$ has the strong epimorphism surjectivity property, it suffices to show that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=A$ for every $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ (see 6.6). To this end, consider $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ and $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. By 4.10 there exist $f\in\mathsf{imp}_{pp}(\mathsf{K})$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$ such that $b=f^{\bm{B}}(a_{1},\dots,a_{n})$. Since $\mathsf{K}$ has the strong Beth definability property, there exists a term $t$ such that $f^{\bm{B}}(a_{1},\dots,a_{n})=t^{\bm{B}}(a_{1},\dots,a_{n})$. Therefore,
@@
This can also be viewed through the lens of 6.7. For let $f$ be the implicit operation of “taking inverses” in monoids and recall that it can be defined by a pp formula (see Theorem 3.7). Clearly, $\mathbb{Z}\leqslant\mathbb{Q}$ and $2\in\mathsf{dom}(f^{\mathbb{Q}})\cap\mathbb{Z}$. Moreover,
@@
Essentially the same argument shows that the quasivariety $\mathsf{RCRing}$ of reduced commutative rings lacks the strong epimorphism surjectivity property. More precisely, let $\mathbb{Z}$ and $\mathbb{Q}$ be the reduced commutative rings of the integers and the rationals, respectively. Moreover, let $f$ be the implicit operation of $\mathsf{RCRing}$ given by 3.18. As $\mathbb{Q}$ is a field, 3.18 guarantees that $f^{\mathbb{Q}}$ is the operation of “taking weak inverses” in $\mathbb{Q}$. Therefore, we can replicate the argument detailed in 6.8, yielding that $\mathsf{RCRing}$ lacks the strong epimorphism surjectivity property, and hence also the strong Beth definability property. ∎

### v2

@@ 4. Existential elimination
As $\bm{A}\leqslant\bm{B}\leqslant\bm{T}_{\mathsf{K}}(\varphi)\in\mathsf{K}$ and $\mathsf{K}$ is a quasivariety with the amalgamation property, we can apply Proposition 4.11 to $y/\theta(\varphi)\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{T}_{\mathsf{K}}(\varphi))\cap B$, obtaining $y/\theta(\varphi)\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. By Theorem 4.10 there exist an $m$-ary $g\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ and $\langle a_{1},\dots,a_{m}\rangle\in\mathsf{dom}(g^{\bm{B}})\cap A^{m}$ such that $g^{\bm{B}}(a_{1},\dots,a_{m})=y/\theta(\varphi)$.
Since $g\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$, there exists a formula $\exists z_{1},\dots,z_{k}\psi(x_{1},\dots,x_{m},y,z_{1},\dots,z_{k})$, where $\psi$ is a conjunction of equations, defining $g$. Together with $\langle a_{1},\dots,a_{m}\rangle\in\mathsf{dom}(g^{\bm{B}})$ and $g^{\bm{B}}(a_{1},\dots,a_{n})=y/\theta(\varphi)$, this guarantees the existence of $b_{1},\dots,b_{k}\in B$ such that
@@
Notice that $\gamma$ is a conjunction of equations because so is $\psi$. Then observe that the formula $\exists z_{1},\dots,z_{k}\psi(x_{1},\dots,x_{m},y,z_{1},\dots,z_{k})$ is functional in $\mathsf{K}$ because it defines $g$. Together with the definition of $\gamma$, this guarantees that $\gamma$ is also functional in $\mathsf{K}$. Hence, $\gamma$ defines some $h\in\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$ by Corollary 3.11. Therefore, to conclude the proof, it suffices to show that $h$ interpolates $f$.
@@
$\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=\{b\in B:$ $\text{ there exist }f\in\mathsf{imp}_{\textsc{eq}}(\mathsf{K})\text{ and }\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$
@@
We close this section with two examples. The first shows that Isbell’s operations cannot be interpolated by any implicit operation defined by a conjunction of equations in the variety of commutative monoids, and the second shows that condition (i) of Theorem 4.3 cannot be improved by requiring the equality $\mathsf{imp}_{\textsc{pp}}(\mathsf{K})=\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$.
@@
For every positive $n$ let $f_{n}$ be the $n$-th Isbell’s operation, viewed as an implicit operation of the variety $\mathsf{Mon}$ of monoids (see Example 3.14 and Theorem 3.15). Then $f_{n}\in\mathsf{imp}_{\textsc{pp}}(\mathsf{Mon})$. We will prove that $f_{n}$ cannot be interpolated by any member of $\mathsf{imp}_{\textsc{eq}}(\mathsf{\mathsf{Mon}})$. Suppose the contrary, with a view to contradiction. Then $f_{n}$ is interpolated by some $g\in\mathsf{imp}_{\textsc{eq}}(\mathsf{Mon})$. Consider the commutative monoids $\mathbb{N}=\langle\mathbb{N};\cdot,1\rangle$ and $\mathbb{Q}=\langle\mathbb{Q};\cdot,1\rangle$. Moreover, let $a_{1},\dots,a_{n},b_{1},\dots,b_{n},c_{1},\dots,c_{2n+1},d$ be the sequence of rationals defined as follows: for every $1\leqslant i\leqslant n$ and $1<j<2n+1$,
@@
Using the formula $\psi_{n}$ in Example 3.14, we have
@@
By the definition of $f_{n}$ this yields $\langle c_{1},\dots,c_{2n+1}\rangle\in\mathsf{dom}(f_{n}^{\mathbb{Q}})$ and $f_{n}^{\mathbb{Q}}(c_{1},\dots,c_{2n+1})=d$. As $g$ interpolates $f_{n}$, we obtain $\langle c_{1},\dots,c_{2n+1}\rangle\in\mathsf{dom}(g^{\mathbb{Q}})$ and $g^{\mathbb{Q}}(c_{1},\dots,c_{2n+1})=d$. Let $\varphi(x_{1},\dots,x_{2n+1},y)$ be the conjunction of equations defining $g$. Then $\mathbb{Q}\vDash\varphi(c_{1},\dots,c_{2n+1},d)$. As $\varphi$ is a universal formula and $c_{1},\dots,c_{2n+1},d\in\mathbb{N}$, Theorem 1.1(iii) implies that $\mathbb{N}\vDash\varphi(c_{1},\dots,c_{2n+1},d)$, and hence $\langle c_{1},\dots,c_{2n+1}\rangle\in\mathsf{dom}(g^{\mathbb{N}})$ and $g^{\mathbb{N}}(c_{1},\dots,c_{2n+1})=d$. Let $\bm{A}$ and $\bm{B}$ be the submonoids of $\mathbb{N}$ with universes $A=\{1\}\cup\{2m:m\in\mathbb{N}\}$ and $B=\{0,1\}$, respectively. Since $c_{1},\dots,c_{2n+1}\in A$ and $g\in\mathsf{imp}_{\textsc{eq}}(\mathsf{Mon})$, Theorem 4.10 yields $3=d\in\mathsf{d}_{\mathsf{Mon}}(\bm{A},\mathbb{N})$. Let $h,k\colon\mathbb{N}\to\bm{B}$ be given by
@@
It is immediate to verify that $h$ and $k$ are homomorphisms such that $h{\upharpoonright}_{A}=k{\upharpoonright}_{A}$ and $h(3)\neq k(3)$, whence $3\notin\mathsf{d}_{\mathsf{Mon}}(\bm{A},\mathbb{N})$. As this is false, we conclude that $f_{n}$ cannot be interpolated by any member of $\mathsf{imp}_{\textsc{eq}}(\mathsf{Mon})$. ∎
@@
We recall that $\mathsf{CCMon}$ has the amalgamation property (see, e.g., [76, pp. 100, 108]) and, therefore, falls within the scope of Theorem 4.3. On the other hand, we will show that $\mathsf{imp}_{\textsc{pp}}(\mathsf{CCMon})\neq\mathsf{imp}_{\textsc{eq}}(\mathsf{CCMon})$.
@@
Notice that $\varphi$ is functional in $\mathsf{CCMon}$. Indeed, $\bm{A}\vDash\varphi(a,b)$ implies $b=1$ for all $\bm{A}\in\mathsf{CCMon}$ and $a,b\in A$. Since $\varphi$ is a pp formula, we can apply Corollary 3.11, obtaining that it defines a unary $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{CCMon})$. Moreover, if $\bm{A}\in\mathsf{CCMon}$, then $\mathsf{dom}(f^{\bm{A}})$ consists of the invertible elements of $\bm{A}$. We show that $f\notin\mathsf{imp}_{\textsc{eq}}(\mathsf{CCMon})$. Suppose the contrary, with a view to contradiction. Then consider the cancellative commutative monoids $\mathbb{N}=\langle\mathbb{N};\cdot,1\rangle$ and $\mathbb{Q}=\langle\mathbb{Q};\cdot,1\rangle$. We have
@@
Since $f$ is defined by a conjunction of equations, from the above display and $1,2\in\mathbb{N}\leqslant\mathbb{Q}$ it follows that $2\in\mathsf{dom}(f^{\mathbb{N}})$, a contradiction with the fact that $2$ is not invertible in $\mathbb{N}$. Hence, we conclude that $\mathsf{imp}_{\textsc{pp}}(\mathsf{CCMon})\neq\mathsf{imp}_{\textsc{eq}}(\mathsf{CCMon})$.
We conclude this example by showing that $\mathsf{imp}_{\textsc{eq}}(\mathsf{CCMon})$ is not closed under composition. Observe that $f$ coincides with the composition $h\circ g$, where $h\in\mathsf{imp}_{\textsc{eq}}(\mathsf{CCMon})$ is the unary implicit operation defined by the equation $y\thickapprox 1$ and $g\in\mathsf{imp}_{\textsc{eq}}(\mathsf{CCMon})$ is the implicit operation of “taking inverses” in monoids (see Example 3.4) restricted to $\mathsf{CCMon}$. Since $h\circ g=f\notin\mathsf{imp}_{\textsc{eq}}(\mathsf{CCMon})$, this shows that $\mathsf{imp}_{\textsc{eq}}(\mathsf{CCMon})$ is not closed under composition. ∎
@@ 5. The strong Beth definability property
The next definition formalizes the idea of interpolating implicit operations by sets of terms and is a particular instance of the notion of interpolation introduced in Definition 4.2.
@@
As a consequence of the Compactness Theorem 1.3, we obtain the following.
@@
Let $f$ be a partial function on an elementary class $\mathsf{K}$. Assume that $f$ is defined by a formula $\varphi(x_{1},\dots,x_{n},y)$ and that it can be interpolated by a set of terms $\{t_{i}:i\in I\}$. Then condition (21) holds. From the Compactness Theorem 1.3 it follows that there exists a finite $T\subseteq\{t_{i}:i\in I\}$ such that
@@
When viewed as an implicit operation on a class of algebras $\mathsf{K}$, every term function $\langle t^{\bm{A}}:\bm{A}\in\mathsf{K}\rangle$ of $\mathsf{K}$ is interpolated by a single term, namely, $t$ (see Example 3.8). However, not all implicit operations can be interpolated by terms. For instance, there is no set of terms interpolating the implicit operation of “taking inverses” in the variety of monoids as we will see in Example 6.8. It is therefore sensible to isolate the cases in which interpolation is always possible, something that indicates a good balance between the expressivity of the language (measured by what can be said in terms of implicit operations) and its actual richness (measured by what can be interpolated, or made explicit, by terms).
@@ 6. The strong epimorphism surjectivity property
We will show that, in the setting of universal classes, the strong Beth definability property is equivalent to the following strengthening of the epimorphism surjectivity property (see Theorem 6.5).
@@
While every class with the strong epimorphism surjectivity property has the epimorphism surjectivity property, the converse need not hold in general. For instance, it is known that only $16$ varieties of Heyting algebras have the strong epimorphism surjectivity property [91, Thm. 8.1]. On the other hand, there exists a continuum of varieties of Heyting algebras with the nonstrong version of this property [14, p. 199]. However, the two properties coincide in quasivarieties with the amalgamation property (see Theorem 7.14).
@@
In the context of quasivarieties, the strong epimorphism surjectivity property admits a purely categorical formulation. For observe that each quasivariety $\mathsf{K}$ can be viewed as a category whose objects are the members of $\mathsf{K}$ and whose arrows are the homomorphisms between them. As quasivarieties contain free algebras (see Theorem 2.19), monomorphisms coincide with embeddings in quasivarieties (see, e.g., [1, Prop. 8.29]). A monomorphism is said to be regular when it is an equalizer. It turns out that a quasivariety $\mathsf{K}$ has the strong epimorphism surjectivity property if and only if all monomorphisms are regular in $\mathsf{K}$ (see, e.g., [76, Prop. 6.1]). ∎
@@
As an immediate consequence of Remark 6.2, we obtain an alternative formulation of the strong epimorphism surjectivity property in terms of dominions.
@@
From Theorem 4.10 and Proposition 6.6 we deduce the following.
@@
A universal class $\mathsf{K}$ has the strong epimorphism surjectivity property if and only if for all $\bm{A}\leqslant\bm{B}\in\mathsf{K}$, $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$, and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$ we have $f^{\bm{B}}(a_{1},\dots,a_{n})\in A$.
We are now ready to prove Theorem 6.5.
@@
Let $\mathsf{K}$ be a universal class. To prove the implication from left to right, we reason by contraposition. Suppose that $\mathsf{K}$ lacks the strong Beth definability property. By Propositions 5.2 and 5.4 there exists $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ that cannot be interpolated by any set of terms. Therefore, there exists $\bm{B}\in\mathsf{K}$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$ for which there exists no term $t$ such that $f^{\bm{B}}(a_{1},\dots,a_{n})=t^{\bm{B}}(a_{1},\dots,a_{n})$. Then $f^{\bm{B}}(a_{1},\dots,a_{n})\notin\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n})$. Let $\bm{A}=\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n})$. Since $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$ and $f^{\bm{B}}(a_{1},\dots a_{n})\notin A$, Corollary 6.7 implies that $\mathsf{K}$ lacks the strong epimorphism surjectivity property.
Next we prove the implication from right to left. Suppose that $\mathsf{K}$ has the strong Beth definability property. In order to prove that $\mathsf{K}$ has the strong epimorphism surjectivity property, it suffices to show that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=A$ for every $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ (see Proposition 6.6). To this end, consider $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ and $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. By Theorem 4.10 there exist $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$ such that $b=f^{\bm{B}}(a_{1},\dots,a_{n})$. Since $\mathsf{K}$ has the strong Beth definability property, there exists a term $t$ such that $f^{\bm{B}}(a_{1},\dots,a_{n})=t^{\bm{B}}(a_{1},\dots,a_{n})$. Therefore,
@@
This can also be viewed through the lens of Corollary 6.7. For let $f$ be the implicit operation of “taking inverses” in monoids and recall that it can be defined by a pp formula (see Theorem 3.7). Clearly, $\mathbb{Z}\leqslant\mathbb{Q}$ and $2\in\mathsf{dom}(f^{\mathbb{Q}})\cap\mathbb{Z}$. Moreover,
@@
Essentially the same argument shows that the quasivariety $\mathsf{RCRing}$ of reduced commutative rings lacks the strong epimorphism surjectivity property. More precisely, let $\mathbb{Z}$ and $\mathbb{Q}$ be the reduced commutative rings of the integers and the rationals, respectively. Moreover, let $f$ be the implicit operation of $\mathsf{RCRing}$ given by Theorem 3.18. As $\mathbb{Q}$ is a field, Theorem 3.18 guarantees that $f^{\mathbb{Q}}$ is the operation of “taking weak inverses” in $\mathbb{Q}$. Therefore, we can replicate the argument detailed in Example 6.8, yielding that $\mathsf{RCRing}$ lacks the strong epimorphism surjectivity property, and hence also the strong Beth definability property. ∎